You have a rigid tank equipped with a pressure gauge. Describe a procedure by which you can use this tank to blend ideal gases in prescribed mole-fraction. partitions.
Answer #1

SOLUTION:

In ideal gases pressures can be taken as the concentrations of the gases. If we have two gases A and B. Lets assume that the mole fraction of each gas is required to be 0.5. We allow gas A to enter the tank and note the pressure exerted by gas A (suppose it is equal to1 bar). Then we will the gas B to enter the tank in such a manner that the partial pressure of B must be equal to partial pressure of A. That is, partial pressure of gas B must be 1 bar as well. Therefore total pressure (partial pressure of A + Partial pressure of B) is equal to 2 bar. Then mole fraction of gas A (xA) is given as:

xA = pA/Ptot = 1/2 = 0.5

similarly

xB = pB/Ptot = 1/2 = 0.5